Rx1nyDxyz:

At a conference for 500 people, 30% of the participants are French, 15% are Americans, 5% are Germans, and are of other nationalities. How many French participants are at the conference? 30 150 160 250

Florisalreadytaken:

what is \[ 500 \times \frac{30}{100} = ? \]

Rx1nyDxyz:

530 over 100?

Rx1nyDxyz:

Oh, sorry. I though you would multiply them across.

Florisalreadytaken:

no -- it'd be \[\frac{ 150\cancel{00}}{ 1\cancel{00} } \]

Florisalreadytaken:

i did it wrong the one be4 this lol -- so what is \[ \frac{150}{1} =? \]

Rx1nyDxyz:

150 i believe.

Florisalreadytaken:

and you're right!

Rx1nyDxyz:

Thank you for helping me with that! :) I really appreciate it!